OpenStudy (anonymous):

tan(x/2-pi/4)=tan(pi/8)

OpenStudy (anonymous):

\[\frac{1}{2}-\frac{1}{4}=\frac{1}{4}\] so i was mistaken this is false

OpenStudy (freckles):

well also tan(x/2-pi/4) varies while tan(pi/8) remains constant